An X-ray beam of 71 pm (1 pm = 10-12m) affects a calcite target.
a) Find the wavelength of the scattered X-rays at an angle of 30 °.
b) What is the largest wavelength change that can be expected in this experiment? (This effect is known as retro dispersion - backscattering)
Answer #1

lamda = 71 pm = 71*10^-12 m

a) the wavelength of the scattered X-rays,

lamda' = lamda + h/(m*c)*(1 - cos(theta))

= 71*10^-12 + 6.625*10^-34/(9.1*10^-31*3*10^8)*(1 - cos(30))

= 7.13*10^-11 m (or) 71.3 pm <<<<<<<-------------Answer

b)

the largest wavelength change that can be expected in this experiment,

lamda' - lamda = h/(m*c)*(1 - cos(180))

= 6.625*10^-34/(9.1*10^-31*3*10^8)*(1 - cos(180))

= 4.85*10^-12 m (or) 4.85 pm <<<<<<<-------------Answer